OpenStudy (anonymous):

y=-3x-6 Find four solutions of the equation. Write the solutions as ordered pairs.

OpenStudy (anonymous):

pick 4 numbers for x (any 4, doesn't matter) then write down the resulting y for the given x so for example if i pick 1 for x, i'd get y=-3*1-6 = -9 (1,-9)

OpenStudy (anonymous):

put x=any value say 1,2,3,4 the find the corresponding value of y

OpenStudy (anonymous):

also: \( \large Since~you~seem~new~here:\\\bf\Huge{~~\color{red}{\boxed{W}}\color{#FF9200}{\boxed{E}}\color{#FFDB00}{\boxed{L}}\color{#B6ff00}{\boxed{C}}\color{#00ff49}{\boxed{O}}\color{#00DBff}{\boxed{M}}\color{#B600ff}{\boxed{E}}\\~~~~~~~~~~~~~~~~~~\boxed{T}\boxed{O}\\~~~~~~~~~~~\color{#0092ff}{\boxed{O}}\color{#0092ff}{\boxed{P}}\color{#0092ff}{\boxed{E}}\color{#0092ff}{\boxed{N}}\color{#7cc517}{\boxed{S}}\color{#7cc517}{\boxed{T}}\color{#7cc517}{\boxed{U}}\color{#7cc517}{\boxed{D}}\color{#7cc517}{\boxed{Y}}\color{#7cc517}{\boxed{!}}} \\ \large\bf You~can~read~a~\underline{\href{ /study#/updates/527023d6e4b08ad7c0d7b5bc }{guide}}~about~openstudy~here.\)

OpenStudy (anonymous):

answers choices are: A.{-4,3} [-7,15] [5,-24] -6,12] B.[-4,6] -7,15] [5,-21] [-6,12] C.[-4,6] [-6,15] [5,-21] [-7,12] D.[4,6] [-7,15] [5,3] [-4,12]

OpenStudy (anonymous):

well, i'd start by plugging in -4 in the formula, and see if the result is either 3 or 6. That way, you eliminate at least 2, and possibly 3 of the answers.